OpenStudy (anonymous):

Write the equation in exponential form. logv4 1/16 = -2

OpenStudy (anonymous):

A. 4 1/2 = 16 B. 4^2 = 16 C. 16 1/2 = 4 D. 4^-2 = 1/16